# Rapid, strongly magnetized accretion in the zero-net-vertical-flux shearing box

### v1

@@ 5.3 The effect of boundary conditions and numerical options
In Fig. 7 we illustrate the effect of changing the vertical boundary conditions and other numerical options on the low-$\beta$ state. As listed in table 1, these simulations explore 4 different vertical boundary conditions based on previous literature (see §3.2), different initial conditions (Eq. 5), a taller box with $L_{z}=15$, and second-order reconstruction with an HLLD Reimman solver, but otherwise have the same parameters (e.g., $c_{s}$). They all reach a quasi-steady low-$\beta$ state and we plot the steady-state time average of $P_{B}=\overline{B_{y}}^{2}$ and $P_{\rm th}=c_{s}^{2}\overline{\rho}$. Overall, we see that the general form is maintained for all cases, although there are certainly differences. Unsurprisingly, given that Simon13 boundary conditions fix the gradients of $B_{x}$ and $B_{y}$ to be large and negative at the boundary, these lead to the steepest $P_{B}$ profile at larger $z$, while the Outflow boundaries, which set $B_{x}^{\prime}=B_{y}^{\prime}=0$, do the opposite (the Lesur13 and Power-Law cases lie between these limits). Less obvious, however, is that these steeper $P_{B}$ profiles lead to states that are more strongly magnetized (a lower-$\beta$ midplane) with larger accretion rates $\alpha$ (see table 1), even though naively one might expect that carrying the flux out of the domain more efficiently would lead to lower midplane magnetic fields. Using $2^{\rm nd}$-order reconstruction leads to the flattest profiles in $P_{B}$ and $\rho$, which appears to be a consequence of it supporting modestly less vigorous turbulence at large $z$, and thus transporting the flux out of the midplane less efficiently (not shown). Finally, $\beta$0.1-tall — with an extended vertical domain (the full extent is not shown), lower resolution per $H_{\rm th}$, and a much lower density floor — produces similar vertical profiles, albeit with lower midplane turbulence levels, which lead to lower $\alpha$ and larger $\beta_{\rm mid}$. This difference is at least partially a consequence of its lower resolution per $H_{\rm th}$ (see App. B), perhaps also influenced by its different domain aspect ratio (see App. A; other than the initial conditions, this case is equivalent to rescaling the vertical box to a lower $H_{\rm th}$, as for $\beta$0.05 or $\beta$0.01).
@@
Various other numerical tests are presented in the Appendices, which describe the dependence on domain aspect ratio (App. A), resolution (App. B), and the density floor (App. C). We briefly summarize these results here. We find that the azimuthal length of the box can have an important influence, with $\beta=0.1$ initial conditions that transition to the low-$\beta$ state in a longer box instead transitioning to the high-$\beta$ state in a shorter box. We interpret this behavior as relating to the Parker instability, which seems to be needed to regenerate the field against vertical escape. The fastest growing Parker-instability modes in this regime are of the “quasi-interchange” variety, which have a specific azimuthal scale (9) that approaches the box scale with shorter boxes, thus potentially interfering with the turbulence. The radial box size does not appear to have an important influence. The dependence on resolution is interestingly non-monotonic, with the transport decreasing with resolution until $1$-$2$ zones per $H_{\rm th}$, then suddenly increasing dramatically at yet lower resolutions. This behavior occurs because the turbulence in the midplane is suppressed with decreasing resolution, thus reducing the transport for similar profiles; but, once the resolution drops so low that it acts to support the collapse of density to the thermal profile in the midplane, the system no longer loses its magnetic flux and supports a $\beta\ll 1$ midplane with higher transport (see §5.6 and e.g., $\beta$0.001 in Fig. 3). Several direct tests show effectively no dependence of the low-$\beta$ state on the choice of density floor.
@@
A final numerical issue is that systems fail to converge to the low-$\beta$ state without a large separation between the vertical domain size and $H_{\rm th}$. Specifically, we have run various test simulations with the same $H_{\rm th}$ as $\beta$1 or $\beta$0.3, but with a stronger initial magnetic field with $\beta<0.1$ (they thus have larger $H_{\beta}$). These cases lose all of their flux and transition into the high-$\beta$ state. This outcome is robust even at resolutions with fewer zones per $H_{\rm th}$ than, e.g., $\beta$R0.1 (showing that it is not due to the stronger turbulent expulsion of flux at higher resolution), and suggests that the low-$\beta$ state can be maintained only when there exists an extended strongly magnetized corona. This could be because the Parker-driven turbulence, as needed to sustain the mean azimuthal field, requires an atmosphere of sufficient vertical extent and/or due to direct effects from the boundary conditions (e.g., the strong vertical flows that develop in their near vicinity).
@@
Overall, we see that while numerical options do have an important impact on the results, as is also the case for the high-$\beta$ shearing-box turbulence (e.g., Guan et al., 2009; Ryan et al., 2017), the general features of the low-$\beta$ state seem robust.
5.4 The necessity of shear and rotation
In Fig. 8 we demonstrate the numerical experiment of re-running two cases that transition into the low-$\beta$ state in the standard shearing box ($\beta$R0.1 and $\beta$0.01), while selectively removing either shear or rotation to diagnose its importance. The top two panels show the effect of removing the mean shear flow; these clearly do not maintain interesting dynamics. The flux continuously escapes in $\beta$R0.1-noshear, while $P_{B}$ becomes flat in $\beta$0.01 due to its Outflow boundary conditions, and both develop perfect thermal profiles in the midplane. The detailed structure at larger $z$ is affected by the density floor and so untrustworthy, but the obvious difference with Fig. 1 nonetheless clearly demonstrates the importance of the shear.
As disussed in §4.1, removing rotation kills the MRI, as well as changing the turbulence properties and Parker instability (see below). The lower panels of Fig. 8 show $\beta$R0.1-norot and $\beta$0.01-norot in which the rotation is artificially removed. Neither of these maintain the low-$\beta$ state, developing a thermally dominated midplane. Interestingly, dynamo cycles develop in $\beta$R0.1-norot, despite the lack of the MRI. Assuming these are of similar origin to those in the high-$\beta$ state (given they have a similar period and structure) it is clear that the nomenclature ‘MRI dynamo cycles’ is not accurate. Given this system’s small $H_{\rm th}$ and the density floor this would be worth a dedicated study in future work. In the lower-$\beta$ case we do not see this feature, presumably due to the poorly resolved midplane and outflow boundary conditions that lead to $P_{B}\approx{\rm const.}$ at $z\gtrsim 2H_{\rm th}$. Notwithstanding, given the clear difference with $\beta$0.01, we can conclude that rotation is crucial for maintaining the low-$\beta$ state.
@@ 5.5 The effect of vertical flux
Figure 9: The effect of a net vertical flux on the low-$\beta$ state. Panels a and b show butterfly diagrams of $\overline{B_{y}}$ (cf. Fig. 1) starting from the same initial conditions as $\beta$R0.1, but with a net vertical flux of $\beta_{z}=1000$ or $\beta_{z}=100$. Diagnostics of the vertical equilibrium (cf. Fig. 3) and turbulence (cf. Fig. 6) are shown in panels d and e, respectively, while panel f shows the contributions to $\alpha$ for the $\beta_{z}=100$ case (cf. Fig. 5). The $\beta_{z}=1000$ field makes little noticeable difference to the morphology, raising $\alpha$ slightly, while $\beta_{z}=100$ causes a stronger mean field, lowering the midplane $\beta$, in line with previous results (Suzuki & Inutsuka, 2009; Bai & Stone, 2013; Salvesen et al., 2016b). Finally, panel c demonstrates the effect of removing the rotation at $\beta_{z}=100$ (cf. Fig. 8), illustrating how the system reverts to high-$\beta$ cycles.
@@
Previous works have found that strongly magnetized ($\beta<1$) profiles are sustained in the presence of a net vertical flux threading the box (e.g., Suzuki & Inutsuka, 2009; Bai & Stone, 2013). We explore here how the ZNVF low-$\beta$ state changes with $\beta_{z}$, finding that it smoothly transitions into the previously explored low-$\beta$ NVF states as $\beta_{z}$ is decreased below $\beta_{z}\simeq 10^{3}$, and that the general shape of the profiles remains rather similar as this occurs. In other words, while for $\beta_{z}\gtrsim 10^{3}$ there exist two qualitatively different self-sustaining accretion states depending on the initial toroidal flux threading the system, for $\beta_{z}\lesssim 10^{3}$ the system always reaches a strongly accreting state characterized by a $\beta\lesssim 1$ midplane, wider density profiles, and a lack of dynamo cycles (much longer-time-duration simulations are shown in Bai & Stone 2013; Salvesen et al. 2016b with Salvesen et al. 2016b’s exhibiting widely spaced, aperiodic flips in polarity at $\beta_{z}=100$). However, in §6 below, we argue that despite the similar appearance of the ZNVF and NVF states, the strong azimuthal field is maintained in a rather different way with strong NVF, effectively being regenerated as a single large-scale MRI channel mode via stretching of the vertical field, as opposed to via a form of turbulent dynamo.
@@
Our results are collected in Fig. 9. In the top panel we show butterfly diagrams for the same setup as $\beta$R0.1 but with $\beta_{z}=10^{3}$ (left) and $\beta_{z}=100$ (middle). The profiles and time evolution appear similar to Bai & Stone (2013); Salvesen et al. (2016b), as well as those with ZNVF (cf. Fig. 1), although with $\beta_{z}=100$ the field becomes significantly stronger and less peaked in the midplane after its initial phase (see also the quasi-steady-state profiles in Fig. 9d). The accretion rate $\alpha$ is also higher at $\beta_{z}=100$ ($\langle{\alpha}\rangle\approx 2.1$ at $\beta_{z}=100$, versus $\langle{\alpha}\rangle\approx 0.7$ and $\langle{\alpha}\rangle\approx 0.6$ for $\beta_{z}=10^{3}$ and $\beta_{z}=\infty$, respectively). [^6] Salvesen et al. (2016b) and Bai & Stone (2013) report $\langle{\alpha}\rangle\simeq 1.0$ at $\beta_{z}=100$ when measured at the midplane or $\langle{\alpha}\rangle\simeq 1.5$ when measured over the full profile. The larger $\langle{\alpha}\rangle$ we report here is due to the larger $L_{z}/H_{\rm th}$ in our simulation, since $\langle{\alpha}\rangle$ is normalized by $\langle{\rho}\rangle$ and $\overline{\rho}$ decreases more rapidly with $z$ than $\overline{B_{x}}\overline{B_{y}}$ (see Fig. 9f). Interestingly, this larger $\langle{\alpha}\rangle$ is not a consequence of the turbulent fluctuations, which are hardly stronger at $\beta_{z}=100$ than at $\beta_{z}=\infty$ (see panel e); instead its accretion stress is nearly flat with $z$ and almost entirely due to the mean field $-\overline{B_{x}}\,\overline{B_{y}}$. This agrees with the conclusions of Bai & Stone 2013 (see their figure 6). The reason relates to how $\overline{B_{x}}$ is generated from $\overline{B_{y}}$ via direct stretching of the net vertical field (as opposed to via the turbulence when $\overline{B_{z}}=0$): this more efficient feedback increases $\alpha$ and lowers $\beta$ due to the stronger fields (see §6). As in previous works, we also observe much stronger outflows at $\beta_{z}\lesssim 10^{3}$, which cause significant mass loss over time if the density is not continuously replenished via an artificial source term. Despite the lower $\beta$ at $\beta_{z}=100$, the density still collapses towards the thermal scale height within $|z|\lesssim H_{\rm th}$, with a similar midplane density structure as the ZNVF runs.
Finally, in Fig. 9c, we illustrate the numerical experiment of removing the rotation in simulations with a strong vertical field $\beta_{z}=100$ (cf. Fig. 8). This causes the system to lose its azimuthal flux, generating cyclic behavior in $\overline{B_{y}}$ and a thermal midplane, rather similar to the high-$\beta$ ZNVF system although with somewhat stronger mean-field cycles. This is consistent with our interpretation that the NVF low-$\beta$ state is effectively a single large-scale MRI mode, which requires rotation to sustain itself, as well as indicating again that the dynamo cycles are not a consequence of the MRI, which does not exist in the absence of rotation.
@@ 5.6 Collapse towards the $\beta\simeq 1$ midplane
Figure 10: Magnetic and thermal pressures (as in Fig. 3) as a function of time, from dark purple to yellow in equal time increments of approximately 1 orbit, over the time range indicated on each panel. Panel (a) shows $\beta$0.1-H4, which starts with $H_{\beta}=4$ and $\beta\approx 0.01$, so that it loses significant flux before reaching a state similar to $\beta$R0.1 (cf. Fig. 7); panel (b) shows $\beta$0.05, which has less flux to lose over its initial phase, but behaves analogously, settling into a similar state (other than its smaller $H_{\rm th}$) once $\beta\simeq 1$ in the midplane. Panel (c) shows $\beta 0.01$-hr, which is refined from the steady state of $\beta 0.01$ at $t_{\rm refine}\approx 23$ (see text). The refined version decreases in time, resettling into a somewhat higher-$\beta$ midplane, even though over the same time period the lower-resolution version $\beta 0.01$ increases its midplane field somewhat (the black lines show $\beta$0.01 at the same time as the yellow lines for $\beta$0.01-hr). This illustrates that the lower-$\beta$ midplanes seen in $\beta$0.01 and $\beta$0.001 (and other cases not shown) are likely the result of poor resolution of $H_{\rm th}$.
Figures 3 & 4 show that even when the disk is initialized with $\beta\ll 1$ (but no vertical field) and reaches the low-$\beta$ state, the midplane is still gas-pressure dominated with $\beta\sim 1$. Here we explore this collapse to a gas pressure-dominated midplane in more detail.
@@
Figure 10 shows several numerical experiments that probe this physics, illustrating the time evolution of $P_{\rm th}$ and $P_{B}$. The top panel shows a simulation ($\beta$0.1-H4) with the same $H_{\rm th}$ and box size as $\beta$R0.1 or $\beta$0.1-mr, but with a more strongly magnetized initial condition with $\beta\approx 0.01$. It thus starts with a wider profile with $H_{\beta}\approx 4$, effectively simulating the small central patch $|z|\lesssim 1.6$ of $\beta$0.01 ($L_{x}$ and $L_{y}$ would also be similarly scaled). After a short reorganization phase from the Guassian initial conditions, the system settles into a quasi-equilibrium state with $P_{B}$ flatter than $P_{\rm th}$ similar to the central region of the generalized power-law profile Eq. 5. This strong field then escapes through the boundary over ${\simeq}15$ orbits, with the density becoming more peaked and $P_{B}$ developing its characteristic “dip” at $z=0$, before halting and reaching steady state at $\beta\sim 1$ (cf. Fig. 7). The same behaviour is observed for $\beta$0.05 (second panel) and other simulations (not shown), albeit with a shorter field-decrease phase because they start from higher $\beta$ initially. In fact, the measured rate of field decrease over this phase, $62\pi\Omega^{-1}d\ln{B_{y}}/dt\approx 1/$ 6 , seems to be independent of boundary conditions, $\beta$, or resolution, as indicated by the thin black line in Fig. 4c. This might be expected for Parker-instability-driven turbulence given its $\beta$-independent growth rate $\gamma\sim\Omega$ (see §4.2). Below, we will argue (§6) that the field is being continuously regenerated over this phase in a similar way to the true saturated state, but that this regeneration is slower than its expulsion through the boundaries.
@@
Two ZNVF simulations in table 1 stand out as exceptions to the above scenario: $\beta$0.01 and $\beta$0.001, which both saturate with a $\beta\ll 1$ midplane (see Fig. 3b; for $\beta$0.001 especially the gas is everywhere magnetically supported, with no $P_{B}$ dip at $z=0$). The thermal scale heights of these simulations are poorly resolved, with only ${\simeq}9$ zones in the thermal midplane ($|z|<H_{\rm th}$) for $\beta$0.01, and only ${\simeq}3$ zones for $\beta$0.001. A natural hypothesis is thus that the resolution halts the collapse of $\rho$ to the thermal scale height thereby artificially halting the collapse at $\beta\ll 1$. We test this in Fig. 10c by taking a snapshot of $\beta$0.01 in the saturated state ($t_{\rm refine}=72\pi/\Omega$) and refining its resolution by a factor of two. [^7] We choose at time at which the shear-periodic boundary conditions are truly periodic in order to refine across the $x$ boundaries. We compare the evolution of this refined simulation ($\beta$0.01-hr) to the unrefined version over the same time period in Fig. 10c. While $P_{B}$ actually increases modestly over this period for $\beta$0.01 (cf. purple and black lines), that of $\beta$0.01-hr decreases (cf. purple and yellow lines). Although it saturates again with $\beta<1$ in $\beta$0.01-hr, this does support the general idea that $\beta\ll 1$ midplanes are likely a result of insufficient resolution, at least in the shearing box and without net vertical flux. Nonetheless, it is interesting — and perhaps physically important despite its unphysical origin — that halting the collapse of the density in the midplane also halts the loss of flux through the boundaries at $z\gg H_{\rm th}$. We will return to this point in the discussion (§7).

### v3

@@ 5.3 The effect of boundary conditions and numerical options
In Fig. 8 we illustrate the effect of changing the vertical boundary conditions and other numerical options on the low-$\beta$ state. As listed in table 1, these simulations explore 4 different vertical boundary conditions based on previous literature (see §3.2), different initial conditions (Eq. 5), a taller box with $L_{z}=15$, and second-order reconstruction with an HLLD Reimman solver, but otherwise have the same parameters (e.g., $c_{s}$). They all reach a quasi-steady low-$\beta$ state and we plot the steady-state time average of $P_{B}=\overline{B_{y}}^{2}$ and $P_{\rm th}=c_{s}^{2}\overline{\rho}$. Overall, we see that the general form is maintained for all cases, although there are certainly differences. Unsurprisingly, given that Simon13 boundary conditions fix the gradients of $B_{x}$ and $B_{y}$ to be large and negative at the boundary, these lead to the steepest $P_{B}$ profile at larger $z$, while the Outflow boundaries, which set $B_{x}^{\prime}=B_{y}^{\prime}=0$, do the opposite (the Lesur13 and Power-Law cases lie between these limits). Less obvious, however, is that these steeper $P_{B}$ profiles lead to states that are more strongly magnetized (a lower-$\beta$ midplane) with larger accretion rates $\alpha$ (see table 1), even though naively one might expect that carrying the flux out of the domain more efficiently would lead to lower midplane magnetic fields. Using $2^{\rm nd}$-order reconstruction leads to the flattest profiles in $P_{B}$ and $\rho$, which appears to be a consequence of it supporting modestly less vigorous turbulence at large $z$, and thus transporting the flux out of the midplane less efficiently. Finally, $\beta$0.1-tall — with an extended vertical domain (the full extent is not shown), fewer grid cells per $H_{\rm th}$, and a much lower density floor — produces similar vertical profiles, albeit with lower midplane turbulence levels, which lead to lower $\alpha$ and larger $\beta_{\rm mid}$. This difference is at least partially a consequence of its coarser resolution per $H_{\rm th}$ (see App. B), perhaps also influenced by its different domain aspect ratio (see App. A; other than the initial conditions, this case is equivalent to rescaling the vertical box to a lower $H_{\rm th}$, as for $\beta$0.05 or $\beta$0.01).
@@
Various other numerical tests are presented in the Appendices, which describe the dependence on domain aspect ratio (App. A), resolution (App. B), and the density floor (App. C). We briefly summarize these results here. We find that the azimuthal length of the box can have an important influence, with $\beta=0.1$ initial conditions that transition to the low-$\beta$ state in a longer box instead transitioning to the high-$\beta$ state in a shorter box. We interpret this behavior as relating to the Parker instability, which seems to be needed to regenerate the field against vertical escape. The fastest growing Parker-instability modes in this regime are of the “quasi-interchange” variety, which have a specific azimuthal scale (9) that approaches the box scale with shorter boxes, thus potentially interfering with the turbulence. The radial box size does not appear to have an important influence. The dependence on resolution is interestingly non-monotonic, with the transport decreasing with coarsening resolution until $1$-$2$ zones per $H_{\rm th}$, then suddenly increasing dramatically at yet coarser resolutions. This behavior arises because the turbulence in the midplane is suppressed with coarser resolution, thus reducing the transport for similar profiles; but, once the resolution drops so low that it acts to support the collapse of density to the thermal profile in the midplane, the system no longer loses its magnetic flux and supports a $\beta\ll 1$ midplane with higher transport (see §5.5 and e.g., $\beta$0.001 in Fig. 4). Several direct tests show effectively no dependence of the low-$\beta$ state on the choice of density floor. Finally, in App. D we demonstrate that both shear and rotation are required to maintain the low-$\beta$ state. Interestingly, in the absence of rotation the system still exhibits high-$\beta$ dynamo cycles; since the MRI does not exist in this case (see §4.1), this suggests that the cyclic dynamo is not associated with the MRI, as usually assumed.
@@
A final numerical issue worth discussing is that the low-$\beta$ state requires a large separation between the vertical domain size and $H_{\rm th}$. This conclusion is based on test simulations at the same $H_{\rm th}$ as $\beta$1 or $\beta$0.3, but with either a stronger initial magnetic field with $\beta<0.1$ (and thus a larger $H_{\beta}$) or with initial radial flux that is sheared into an azimuthal field in the simulation’s early stages (as in e.g., $\beta$R0.1). These tests lose their flux and transition into the high-$\beta$ state. The same behavior occurs at resolutions with fewer zones per $H_{\rm th}$ than some low-$\beta$ simulations (e.g., $\beta$R0.1), showing that it is not an indirect consequence of changing the resolution per $H_{\rm th}$. This suggests that the low-$\beta$ state can be maintained only when there exists an extended strongly magnetized corona, perhaps because the Parker-driven dynamo requires an atmosphere of sufficient vertical extent and/or due to direct effects from the boundary conditions.
@@
Overall, we see that while numerical options do have an important impact on the results, as is also the case for the high-$\beta$ shearing-box turbulence (e.g., Guan et al., 2009; Gressel, 2013; Ryan et al., 2017), the general features of the low-$\beta$ state seem robust.
@@ 5.4 The effect of vertical flux
Figure 9: The effect of a net vertical flux on the low-$\beta$ state. Panels a and b show butterfly diagrams of $\overline{B_{y}}$ (cf. Fig. 2) starting from the same initial conditions as $\beta$R0.1, but with a net vertical flux of $\beta_{z}=1000$ or $\beta_{z}=100$. Diagnostics of the vertical equilibrium (cf. Fig. 4) and turbulence (cf. Fig. 7) are shown in panels c and d, respectively, while panel e shows the contributions to $\alpha(z)$ for the $\beta_{z}=100$ case (cf. Fig. 6). The $\beta_{z}=1000$ field makes little noticeable difference to the morphology, raising $\alpha$ slightly, while $\beta_{z}=100$ causes a stronger mean field, lowering the midplane $\beta$, in line with previous results (Suzuki & Inutsuka, 2009; Bai & Stone, 2013; Salvesen et al., 2016b).
@@
Previous works have found that strongly magnetized ($\beta<1$) profiles are sustained in the presence of a net vertical flux threading the box (e.g., Suzuki & Inutsuka, 2009; Bai & Stone, 2013). We explore here how the ZNVF low-$\beta$ state changes with $\beta_{z}$, finding that it smoothly transitions into this low-$\beta$ NVF state as $\beta_{z}$ decreases below $\beta_{z}\simeq 10^{3}$. The general shape of the profiles remains rather similar as this occurs. In other words, while for $\beta_{z}\gtrsim 10^{3}$ there exist two qualitatively different self-sustaining accretion states depending on the initial toroidal flux threading the system, for $\beta_{z}\lesssim 10^{3}$ the system always reaches a strongly accreting state characterized by a $\beta\lesssim 1$ midplane, wider density profiles, and a lack of dynamo cycles (much longer-time-duration simulations are shown in Bai & Stone 2013; Salvesen et al. 2016b with Salvesen et al. 2016b’s exhibiting widely spaced, aperiodic flips in polarity at $\beta_{z}=100$). However, as argued in §6 below, despite the similar appearance of the ZNVF and NVF states, the mechanism that sustains the strong azimuthal field with NVF is very different: NVF drives a single large-scale MRI channel mode, while the ZNVF state is supported by the turbulent dynamo.
@@
Our results are collected in Fig. 9. In the top panel we show butterfly diagrams for the same setup as $\beta$R0.1 but with $\beta_{z}=10^{3}$ (panel a) and $\beta_{z}=100$ (panel b). The profiles and time evolution appear similar to Bai & Stone (2013); Salvesen et al. (2016b), as well as those with ZNVF (cf. Fig. 2), although with $\beta_{z}=100$ the field becomes significantly stronger and less peaked in the midplane after its initial phase (see also the quasi-steady-state profiles in Fig. 9c). The accretion rate is also higher at $\beta_{z}=100$ (${\alpha}\approx 2.1$ at $\beta_{z}=100$, versus ${\alpha}\approx 0.7$ and ${\alpha}\approx 0.6$ for $\beta_{z}=10^{3}$ and $\beta_{z}=\infty$, respectively). [^8] Salvesen et al. (2016b) and Bai & Stone (2013) report ${\alpha}\simeq 1.0$ at $\beta_{z}=100$ when measured at the midplane or ${\alpha}\simeq 1.5$ when measured over the full profile. The larger ${\alpha}$ we report here is due to the larger $L_{z}/H_{\rm th}$ in our simulation, since ${\alpha}$ is normalized by $\langle{\rho}\rangle$ and $\overline{\rho}$ decreases more rapidly with $z$ than $\overline{B_{x}}\,\overline{B_{y}}$ (see Fig. 9e). This larger ${\alpha}$ at lower $\beta_{z}$ is not a consequence of the turbulent fluctuations, which are hardly stronger at $\beta_{z}=100$ than at $\beta_{z}=\infty$ (see panel d); instead its accretion stress is nearly flat with $z$ and almost entirely due to the mean field $-\overline{B_{x}}\,\overline{B_{y}}$ (see panel e). This agrees with the conclusions of Bai & Stone 2013 (see their figure 6). This behavior results from $\overline{B_{x}}$ being generated from $\overline{B_{y}}$ via direct stretching of the net vertical field (as opposed to via the turbulence in ZNVF simulations): this more efficient feedback, which is effectively the aforementioned domain-scale MRI channel mode, increases $\alpha$ and lowers $\beta$ due to the stronger fields (see §6). As in previous works, we also observe much stronger outflows at $\beta_{z}\lesssim 10^{3}$, which cause significant mass loss over time if the density is not continuously replenished via the artificial source term. Despite the lower $\beta$ at $\beta_{z}=100$, the density still collapses towards the thermal scale height within $|z|\lesssim H_{\rm th}$, with a similar midplane density structure as the ZNVF runs.
@@ 5.5 Collapse towards the $\beta\simeq 1$ midplane
Figure 10: Magnetic and thermal pressures (as in Fig. 4) as a function of time, from dark purple to yellow in equal time increments of approximately 1 orbit, over the time range indicated on each panel. The top panel shows $\beta$0.1-H4, which starts with $H_{\beta}=4$ and $\beta\approx 0.01$, so that it loses significant flux before reaching a state similar to $\beta$R0.1 (cf. Fig. 8); the middle panel shows $\beta$0.05, which has less flux to lose over its initial phase, but behaves analogously, settling into a similar state (other than its smaller $H_{\rm th}$) once $\beta\simeq 1$ in the midplane. The lower panel shows $\beta 0.01$-hr, which is refined from the steady state of $\beta 0.01$ at $t_{\rm refine}\approx 72\pi/\Omega$ (see text). The refined version decreases in time, resettling into a somewhat higher-$\beta$ midplane, even though over the same time period the coarser-resolution version $\beta 0.01$ increases its midplane field somewhat (the black lines show $\beta$0.01 at the same time as the yellow lines for $\beta$0.01-hr). This illustrates that the lower-$\beta$ midplanes seen in $\beta$0.01 and $\beta$0.001 (and other cases not shown) are likely the result of poor resolution of $H_{\rm th}$.
Figures 4 & 5 show that even when ZNVF simulations are initialized with $\beta\ll 1$ and reach the low-$\beta$ state, the midplane usually reaches a gas-pressure dominated $\beta\sim 1$ steady state. Here we explore how the system behaves as it loses magnetic flux and evolves towards a $\beta\sim 1$ midplane, also showing that the process is sensitive to resolving the thermal scale height.
@@
Figure 10 shows several numerical experiments that probe this physics, illustrating the time evolution of $P_{\rm th}$ and $P_{B}$. The top panel shows a simulation ($\beta$0.1-H4) with the same $H_{\rm th}$ and box size as $\beta$R0.1 or $\beta$0.1-mr, but with a more strongly magnetized initial condition with $\beta\approx 0.01$. It thus starts with a wider profile with $H_{\beta}\approx 4$, effectively simulating the small central patch $|z|\lesssim 1.6$ of $\beta$0.01 ($L_{x}$ and $L_{y}$ would also be similarly scaled). After a short reorganization phase from the Guassian initial conditions, the system settles into a quasi-equilibrium state with $P_{B}$ flatter than $P_{\rm th}$, similar to the central region of the generalized power-law profile Eq. 5 (see Fig. 1). This strong field then escapes through the boundary over ${\simeq}15$ orbits, with the density becoming more peaked and $P_{B}$ developing its characteristic “dip” at $z=0$, before halting and reaching steady state at $\beta\sim 1$ (cf. Fig. 8). The same behaviour is observed for $\beta$0.05 (middle panel) and other simulations (not shown), albeit with a shorter field-decrease phase because they start from higher $\beta$ initially. In fact, the measured rate of field decrease over this phase, $62\pi\Omega^{-1}d\ln{B_{y}}/dt\approx 1/$ 6 , seems to be independent of boundary conditions, $\beta$, or resolution, as indicated by the thin black line in the lower panel of Fig. 5. This might be expected for Parker-instability-driven turbulence given its $\beta$-independent growth rate $\gamma\sim\Omega$ (see §4.2). Below (§6), we will argue that the field is being continuously regenerated over this phase in a similar way to the true saturated state, but that this regeneration is slower than its expulsion through the boundaries.
@@
Two ZNVF simulations in table 1 stand out as exceptions to the above scenario: $\beta$0.01 and $\beta$0.001, which both saturate with a $\beta\ll 1$ midplane (see Fig. 4b; for $\beta$0.001 especially the gas is everywhere magnetically supported, with no $P_{B}$ dip at $z=0$). The thermal scale heights of these simulations are poorly resolved, with only ${\simeq}9$ zones in the thermal midplane ($|z|<H_{\rm th}$) for $\beta$0.01, and only ${\simeq}3$ zones for $\beta$0.001. A natural hypothesis is thus that the resolution halts the collapse of $\rho$ to the thermal scale height thereby artificially halting the collapse at $\beta\ll 1$. We test this in the lower panel of Fig. 10 by taking a snapshot of $\beta$0.01 in the saturated state ($t_{\rm refine}=72\pi/\Omega$) and refining its resolution by a factor of two. [^9] At $t_{\rm refine}$ the shear-periodic boundary conditions are truly periodic, allowing straightforward refinement over the $x$ boundaries. We compare the evolution of this refined simulation ($\beta$0.01-hr) to the unrefined version over the same time period. While $P_{B}$ actually increases modestly over this period for $\beta$0.01 (cf. purple and black lines), that of $\beta$0.01-hr decreases (cf. purple and yellow lines). Although it saturates again with $\beta<1$ in $\beta$0.01-hr, this does support the general idea that $\beta\ll 1$ midplanes are likely a result of insufficient resolution, at least in the shearing box and without net vertical flux. Nonetheless, it is interesting — and perhaps physically important despite its unphysical origin — that halting the collapse of the density in the midplane also halts the loss of flux through the boundaries at $z\gg H_{\rm th}$. We will return to this point in the discussion (§7).
@@
Figure 12: Terms in the equation for $\partial_{t}\overline{B_{y}}$ in $\beta$R0.1, averaged over the steady state in time (see Eq. 10). We see that near the midplane, $B_{y}$ flux is supplied by the shearing of $\overline{B_{x}}$, which is balanced by transport to higher $|z|$ from the turbulence ($\partial_{z}\mathcal{E}_{x}$) with small contributions from mean outflows on either side of the midplane. In contrast, at large $|z|$, $\overline{B_{y}}$ is sustained primarily by the upwards transport of flux by the turbulence ($\partial_{z}\mathcal{E}_{x}$), which competes against strong vertical outflows through the boundary. This shows that, absent a dynamo effect to regenerate $\overline{B_{x}}$, the field would decay within several orbits. The thin purple line compares the measured $\partial_{z}\mathcal{E}_{x}$ with the dynamo-theory result $\mathcal{E}_{x}\approx\eta_{\rm turb}\partial_{z}B_{y}$ with $\eta_{\rm turb}=\Omega^{-1}\overline{\delta\bm{u}^{2}}/3$ (we smooth this with a 5-gridpoint moving average for clarity).